# A Catalog of $\sim 146,000$ Radio Galaxies up to $z\sim 5$

### v1

@@ 3.5 Cross-matches between Surveys
We search for LoTSS 144 MHz radio counterparts for the UNIONS-VLASS radio galaxies by applying a positional cross-match of $<2.5\arcsec$ radius, consistent with the VLASS beam size. This cross-match results in 101,671 radio galaxies identified in UNIONS, VLASS, and LoTSS. Finally, there are 64,672 UNIONS-VLASS-FIRST-LoTSS radio galaxies.
@@ 3.6.1 Valid Photometric Redshifts of UNIONS
The photometric redshifts presented in UNIONS are estimated from Bayesian Photometric Redshift (BPZ; Benítez, 2000). The source with reliable photometric redshift should have valid photometric measurements in all five bands ($ugriz$): (1) the photometry is not flagged (column $\mathtt{FLAG\_GAaP\_{ugriz}==0}$) and (2) the AB magnitudes within the optimal Gaussianized aperture should be valid measurements (column $\mathtt{MAG\_GAaP\_{ugriz}!=(-99|99})$), where $\mathtt{99}$ indicates that the source is below the detection limit and $\mathtt{-99}$ stands for an absence of the observational data. Therefore, of 146,212 “clean” UNIONS-VLASS sources, 20,358 have photo-$z$ with good reliabilities. These sources are labeled with $\mathtt{0}$ in the column $\mathtt{flag\_z\_photo}$.
[…]
@@ 3.6.2 Photometric versus Spectroscopic Redshifts
Gwyn et al. (2025) compared ${z}_{\mathrm{photo}}$ estimates with the spectroscopic redshifts collected from publicly available datasets in the UNIONS footprint. The bias and scatter of photo-$z$ are represented by the median and normalized median absolute deviation (NMAD) of the quantity $\frac{\Delta z}{1+z}=\frac{{z}_{\mathrm{spec}}-{z}_{\mathrm{photo}}}{1+{z}_{\mathrm{spec}}}$, respectively. Correspondingly, the bias is $-0.008$ and the scatter is $0.061$ for UNIONS sources with ${z}_{\mathrm{photo}}\lesssim 1.6$. If selecting outliers by $\frac{|\Delta z|}{1+z}>0.15$, then $16.3\%$ of the sources are considered having no reliable photo-$z$ estimates.
Figure 6: Comparisons between UNIONS photometric and SDSS spectroscopic redshifts for ${\color[rgb]{0,0,0}2,175}$ UNIONS-VLASS sources. The solid line indicates ${z}_{\mathrm{photo}}={z}_{\mathrm{spec}}$ while the dashed lines indicate $({z}_{\mathrm{photo}}-{z}_{\mathrm{spec}})/(1+{z}_{\mathrm{spec}})=\pm 0.1$, which represents the photo-$z$ scatter for guidance.
[…]
@@ 3.6.3 Extended Radio Sources
The morphology of radio emission, such as its compactness, presumably originating from radio jets, provides key information about the evolution of RLAGNs. To distinguish between compact (unresolved) and extended (resolved) sources, we make use of the ratio of the total radio flux density (${F}_{\mathrm{int}}$) to the peak radio flux density (${F}_{\mathrm{peak}}$), a widely adopted method (e.g. Ivezić et al., 2002; Bondi et al., 2003; Schinnerer et al., 2007). This ratio directly measures the extension of a radio source, and for unresolved or marginally resolved sources, ${F}_{\mathrm{int}}={F}_{\mathrm{peak}}$ (e.g., Bondi et al., 2008). Extended and resolved sources have ${F}_{\mathrm{int}}>{F}_{\mathrm{peak}}$ whereas the background noise could lower the total flux density. Therefore, we plot ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}$ as a function of the signal-to-noise ratio $\mathrm{S/N}\ (={F}_{\mathrm{peak}}/\mathrm{rms})$ in Figure 7 for 146,212 “clean” UNIONS-VLASS sources (blue crosses) and those (79,638) with FIRST detections (orange dots). Values of ${F}_{\mathrm{int}}<{F}_{\mathrm{peak}}$ then a result of the measurement uncertainties probably arising from the background noises that could lower ${F}_{\mathrm{int}}$ (e.g., Bondi et al., 2003).
To select extended sources, we first determined a lower envelope that encloses 99% of the UNIONS-VLASS sources with ${F}_{\mathrm{int}}<{F}_{\mathrm{peak}}$, as indicated by the dashed cyan curve. We then mirrored it above the horizontal line defined by ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1$, and this upper envelope (solid cyan curve) is represented as:
@@
Adopting these criteria, 29,862 ($\sim 20\%$) of our UNIONS-VLASS sources are extended. We performed a similar evaluation for sources detected in FIRST where the FIRST S/N is defined as $({F}_{\mathrm{peak}}-0.25)/\rm{rms}$ to correct for the CLEAN bias (Helfand et al., 2015), and the corresponding upper envelope is defined by ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+7.9\times\left[\left({{F}_{\mathrm{peak}}-0.25}\right)/\mathrm{rms}\right]^{-1}$, which is in general agreement with ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+6.5\times\left({{F}_{\mathrm{peak}}/\mathrm{rms}}\right)^{-1}$ calculated for the HSC-FIRST catalog (Yamashita et al., 2018).
Alternatively, thanks to the higher sensitivity and better resolution, the VLA-COSMOS 3 GHz survey (Smolčić et al., 2017) determined the envelope by a 95% enclosure of the sources below the ${F}_{\mathrm{int}}={F}_{\mathrm{peak}}$ line. Adopting this criterion, the upper envelope is
@@ 3.6.4 Additional Columns
We calculate the angular separations between each pair of UNIONS, VLASS, FIRST, and LoTSS sources, and they are recorded in the corresponding columns starting with $\mathtt{sep\_\{survey1\}\_\{survey2\}}$. We calculate the spectral indices between each pair of radio surveys if there are detections, and the corresponding columns start with $\mathtt{alpha\_\{frequency1\}\_\{frequency2\}}$. Since the spectral index is sensitive to uncertainties in the flux densities, we provide the corresponding uncertainties in the spectral index in the columns $\mathtt{alpha\_\{frequency1\}\_\{frequency2\}\_err}$.
We calculate the spectral luminosities [$\mathrm{\ W\ Hz^{-1}}$] at 1.4 and 3 GHz using the following equation:
@@
[…]
  • detected in FIRST but not in LoTSS: $S_{\mathrm{obs}}=S_{\mathrm{FIRST}}$ and ${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}=-0.7$ used (Gasperin et al., 2018);
@@
The spectral luminosities will be discussed in detail in § 4.3. The calculations and discussions of radio AB magnitudes ($\mathtt{mag\_\{survey\}}$) and radio loudness ($\mathtt{r\_obs}$) will be presented in § 4.1 and § 4.3, respectively. A full list and details of the columns are given in Table LABEL:tab:_catalog_columns.
Figure 9: Left: VLASS 3 GHz source counts as a function of the AB radio magnitude t. Source counts at $15.511.5<t<$ 15.5 are fitted with a linear function, and the best fits are indicated by the colored line, while the solid ones are fittings in this work and dashed ones are from the literature. The black circles represent UNIONS-VLASS sources, black squares represent the “all” VLASS Epoch 2 catalog with criteria (except $\mathtt{NN\_dist}>30\arcsec$) in § 2.2 applied, and golden crosses represent UNIONS-VLASS sources detected in FIRST, and the size of each sample is shown in the parenthesis. Poisson errors in the sampling bins are represented by vertical bars. Right: Radio number counts normalized to the static Euclidean geometry.
@@ 4 Radio and Optical Number Counts
We here summarize the number counts and basic radio properties of AGNs in the UNIONS-VLASS catalog, with the comparisons from previous radio galaxy catalogs in the literature.
@@ 4.1 UNIONS-VLASS 3 GHz Radio Source Counts
Radio source number counts are often normalized to Euclidean geometry and examined on the plane of flux density ($S$) and $S^{2.5}\ dN/dS$, where $N$ is the number of sources within each flux density bin (right panel of Figure 9). In this work, we use the plane of the radio AB magnitude t and the number counts N in each magnitude bin to make direct comparisons with other catalogs at 1.4 and 3 GHz. Here, the radio AB magnitude is defined as:
$t\ [\mathrm{AB\ mag}]=-2.5\log\left({{F}_{\mathrm{int}}/3631\ \mathrm{Jy}}\right).$ (4)
The $3\sigma\ (=0.36$ mJy) detection limit of VLASS then corresponds to 17.5 and the $5\sigma\ (=1$ mJy) detection limit of FIRST then corresponds to 16.4.
@@
In the left panel of Figure 9, we show radio source counts for 146,212 “clean” UNIONS-VLASS radio galaxies (black circles). We make a UNIONS-VLASS subsample for 79,638 sources that are also detected in FIRST (golden crosses) and compare it with 1.4 GHz radio galaxy catalogs employing FIRST (Ivezić et al., 2002; Yamashita et al., 2018). We also apply the criteria described in § 2.2, except $\mathtt{NN\_dist>30\arcsec}$, to the entire VLASS2 catalog, resulting in 1,873,519 sources (black squares). We compare this “all” VLASS2 sample with the “clean” VLASS2 sample that contains 1,380,707 sources used for the cross-match. The distributions of these samples at $15.511.5<t<$ 15.5 are fitted with a linear function and are indicated by the solid lines. The best fits of the data collected from the literature are indicated by the dashed lines and are summarized in Table 2.
@@
The trends between “all” VLASS (blue solid line) and UNIONS-VLASS sources (black solid line) are in rough consistency. The global difference and the larger discrepancy at the bright end are attributed to the fact that we have applied $\mathtt{NN\_dist>30\arcsec}$ to VLASS2 in the pre-cross-match stage. This criterion may remove a significant fraction of low-$z$ sources that have bright radio lobes and hotspots identified as adjacent components, contributing to the number counts at the bright end. If we compare UNIONS-VLASS with the “clean” VLASS2 used for the cross-match, then this discrepancy diminishes, and the global difference is consistent with the matching rate of $0.594$.
@@
Compared with 1.4 GHz HSC- and SDSS-FIRST source counts (golden and cyan dashed lines), our 3 GHz UNIONS-VLASS sources (black circles) have a slightly steeper increase at $15.511.5<t<$ 15.5. This could be explained by our choice of $\mathtt{NN\_dist>30\arcsec}$. The best fit of the subsample of our UNIONS-VLASS sources with FIRST detections (golden solid line) has a slope in general consistency with previous studies. This comes from the fact that the sky coverage of UNIONS-VLASS is beyond the survey area of FIRST, thus resulting in an underestimate of the real number counts for our UNIONS-VLASS-FIRST sample.
@@
All VLASS $-3.509\ (\pm 0.269)$ $0.317\ (\pm 0.018)$
UNIONS-VLASS $-4.870\ (\pm 0.067)$ $0.391\ (\pm 0.005)$
UNIONS-VLASS-FIRST $-4.779\ (\pm 0.142)$ $0.371\ (\pm 0.010)$
$r-$band $-0.932\ (\pm 0.067)$ $0.074\ (\pm 0.003)$

### v2

@@ 3.5 Cross-matches between Surveys
We search for LoTSS 144 MHz radio counterparts for the UNIONS-VLASS radio galaxies by applying a positional cross-match of $2.5\arcsec$ radius, consistent with the VLASS beam size. This cross-match results in 101,671 radio galaxies identified in UNIONS, VLASS, and LoTSS. Finally, there are 64,672 UNIONS-VLASS-FIRST-LoTSS radio galaxies.
@@ 3.6.1 Valid Photometric Redshifts of UNIONS
The photometric redshifts presented in UNIONS are estimated from Bayesian Photometric Redshift (BPZ; benitez_bayesian_2000). The source with reliable photometric redshift should have valid photometric measurements in all five bands (ugrizugriz): (1) the photometry is not flagged (column $\mathtt{FLAG\_GAaP\_{ugriz}==0}$) and (2) the AB magnitudes within the optimal Gaussianized aperture should be valid measurements (column $\mathtt{MAG\_GAaP\_{ugriz}!=(-99|99})$), where $\mathtt{99}$ indicates that the source is below the detection limit and $\mathtt{-99}$ stands for an absence of the observational data. Therefore, of 146,212 “clean” UNIONS-VLASS sources, 20,358 have photo-$z$ with good reliabilities. These sources are labeled with $\mathtt{0}$ in the column $\mathtt{flag\_z\_photo}$.
To maximize the availability of the redshift information in our catalog, we further make use of the 4-band version of UNIONS5000 ($\mathrm{UNIONS5000}_{ugri}$ hereafter) in which the photometric redshifts are estimated based on ugriugri, without the $z$-band photometry. We perform an angular separation-based cross-match between VLASS2 and $\mathrm{UNIONS5000}_{ugri}$ and exclude the cases where multiple optical sources are associated with one radio source. This is to circumvent erroneously associating the sources in $\mathrm{UNIONS5000}_{ugri}$ with another one in $\mathrm{UNIONS5000}_{ugriz}$. This cross-matching yields $\color[rgb]{0,0,0}118,370$} radio galaxies whose radio sources have a unique optical counterpart within the search radius. Adopting the criteria $\mathtt{FLAG\_GAaP\_{ugri}==0}$ and $\mathtt{MAG\_GAaP\_{ugri}!=(-99|99})$, $\color[rgb]{0,0,0}59,317$} have valid redshifts. Excluding those with valid photo-$z$ in $\mathrm{UNIONS5000}_{ugriz}$ as well, this adds up $\color[rgb]{0,0,0}42,661$} photo-$z$ estimates to our catalog (indicated by $\mathtt{flag\_z\_photo==1}$). And in total, and there are $\color[rgb]{0,0,0}63,019$} reliable ${z}_{\mathrm{photo}}$.
@@ 3.6.2 Photometric versus Spectroscopic Redshifts
gwyn_unions_2025 compared ${z}_{\mathrm{photo}}$ estimates with the spectroscopic redshifts collected from publicly available datasets in the UNIONS footprint. The bias and scatter of photo-$z$ are represented by the median and normalized median absolute deviation (NMAD) of the quantity $\frac{\Delta z}{1+z}=\frac{{z}_{\mathrm{spec}}-{z}_{\mathrm{photo}}}{1+{z}_{\mathrm{spec}}}$, respectively. Correspondingly, the bias is 0.008-0.008 and the scatter is $0.061$ for UNIONS sources with ${z}_{\mathrm{photo}}\lesssim 1.6$. If selecting outliers by $\frac{|\Delta z|}{1+z}>0.15$, then $16.3\%$ of the sources are considered having no reliable photo-$z$ estimates.
Figure 6: Comparisons between UNIONS photometric and SDSS spectroscopic redshifts for $\color[rgb]{0,0,0}12,798$} UNIONS-VLASS sources. The solid line indicates ${z}_{\mathrm{photo}}={z}_{\mathrm{spec}}$ while the dashed lines indicate $({z}_{\mathrm{photo}}-{z}_{\mathrm{spec}})/(1+{z}_{\mathrm{spec}})=\pm 0.1$, which represents the photo-$z$ scatter for guidance.
In Figure 6, we compare the valid ${z}_{\mathrm{photo}}$ of our UNIONS-VLASS sources with the ${z}_{\mathrm{spec}}$ collected from SDSS DR16 (ahumada_16th_2020) to check the reliability of the photo-zzestimates, and $\color[rgb]{0,0,0}12,798$} sources are included in this comparison. The bias of the photo-$z$ estimates of these sources is $\color[rgb]{0,0,0}\sim 0.002$}, and this bias remains for those at ${z}_{\mathrm{photo}}<1$. We calculate the MAD of $\frac{\Delta z}{1+z}$ and define outliers according to the NMAD. For all with ${z}_{\mathrm{spec}}$, the MAD is $0.038$, corresponding to bias $\color[rgb]{0,0,0}\sim 0.002$}, and correspondingly, the NMAD is $\color[rgb]{0,0,0}1.4826\times 0.038\approx 0.057$}. For those at ${z}_{\mathrm{photo}}>1$, the bias worsens to $\color[rgb]{0,0,0}0.026$} and the scatter worsens to $\color[rgb]{0,0,0}0.213$}. Adopting the criterion of $\frac{\Delta z}{1+z}>\mathrm{NMAD}\ (=0.057)$, $\color[rgb]{0,0,0}5264\ (41.1\%)$} are outliers. If choosing $\frac{\Delta z}{1+z}>0.15$ instead, 3540 (27.7%) are outliers, which is still higher than that of UNIONS (gwyn_unions_2025). This large fraction of outliers is likely attributed to sources whose spectroscopic redshifts are much larger than photometric ones, namely, $|{z}_{\mathrm{spec}}-{z}_{\mathrm{photo}}|>0.5$, occupying 23.3% ($\color[rgb]{0,0,0}2981$}) of the total sample. These large discrepancies are much more severe at ${z}_{\mathrm{photo}}<1.5$. As we will discuss in § 5.2, in which we select high-$z$ RLAGNs via the dropout-selection method, this could suggest an underestimate of the number density of high-$z$ sources in UNIONS. This systematic underestimation in redshift is probably due to the absence of AGN-dominated templates in the photo-z SED fitting procedure used by UNIONS (gwyn_unions_2025). As a consequence, these RGs with large discrepancies between ${z}_{\mathrm{photo}}$ and ${z}_{\mathrm{spec}}$ are poor fits by the galaxy templates, because radio-selected sources can exhibit diverse optical properties – ranging from low-luminosity radio AGNs in the nearby Universe to high-$z$ luminous RL quasars that outshine their host galaxies (duncan_all-purpose_2022). Nonetheless, we see good agreement between ${z}_{\mathrm{spec}}$ and ${z}_{\mathrm{photo}}$ at $z\gtrsim 2.5$. Therefore, our UNIONS-VLASS catalog provides a promising exploration for high-z RLAGNs.
@@ 3.6.3 Extended Radio Sources
The morphology of radio emission, such as its compactness, presumably originating from radio jets, provides key information about the evolution of RLAGNs. To distinguish between compact (unresolved) and extended (resolved) sources, we make use of the ratio of the total radio flux density (${F}_{\mathrm{int}}$) to the peak radio flux density (${F}_{\mathrm{peak}}$), a widely adopted method (e.g. ivezic_optical_2002; bondi_vla-virmos_2003; schinnerer_vla-cosmos_2007). This ratio directly measures the extension of a radio source, and for unresolved or marginally resolved sources, ${F}_{\mathrm{int}}={F}_{\mathrm{peak}}$ (e.g., bondi_vla-cosmos_2008). Extended and resolved sources have ${F}_{\mathrm{int}}>{F}_{\mathrm{peak}}$ whereas the background noise could lower the total flux density. Therefore, we plot ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}$ as a function of the signal-to-noise ratio $\mathrm{S/N}\ (={F}_{\mathrm{peak}}/\mathrm{rms})$ in Figure 7 for 146,212 “clean” UNIONS-VLASS sources (blue crosses) and those (79,638) with FIRST detections (orange dots). Values of ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}<1$ are then attributed to the measurement uncertainties probably arising from the background noises that could lower ${F}_{\mathrm{int}}$ (e.g., bondi_vla-virmos_2003).
To select extended sources, we first determine a lower envelope that encloses 99% of the UNIONS-VLASS sources with ${F}_{\mathrm{int}}<{F}_{\mathrm{peak}}$, as indicated by the dashed cyan curve. We then mirror it above the horizontal line defined by ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1$, and this upper envelope (solid cyan curve) is represented as:
@@
Adopting these criteria, 29,862 ($\sim 20\%$) of our UNIONS-VLASS sources are extended. We perform a similar evaluation for sources detected in FIRST, where the FIRST S/N is defined as $({F}_{\mathrm{peak}}-0.25)/\rm{rms}$ to account for the CLEAN bias (helfand_last_2015). The corresponding upper envelope is given by ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+7.9\times\left[\left({{F}_{\mathrm{peak}}-0.25}\right)/\mathrm{rms}\right]^{-1}$, which is generally in line with ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+6.5\times\left({{F}_{\mathrm{peak}}/\mathrm{rms}}\right)^{-1}$ derived for the HSC-FIRST catalog (yamashita_wide_2018).
Alternatively, thanks to the higher sensitivity and better resolution, the VLA-COSMOS 3 GHz survey (smolcic_vla-cosmos_2017) determined the envelope by a 95% enclosure of the sources below the ${F}_{\mathrm{int}}={F}_{\mathrm{peak}}$ line. Adopting this criterion, the upper envelope is
@@ 3.6.4 Additional Columns
In the final UNIONS-VLASS radio galaxy catalog, we provide $\mathtt{Z\_B\_ref}$ for the preferred UNIONS photo-$z$. If $\mathtt{flag\_z\_photo==0}$, ${z}_{\mathrm{photo}}$ of $\mathrm{UNIONS5000}_{ugriz}$ is used; if $\mathtt{flag\_z\_photo==1}$, ${z}_{\mathrm{photo}}$ of $\mathrm{UNIONS5000}_{ugri}$ is used; and if $\mathtt{flag\_z\_photo==-1}$, the photometric redshifts are those from $\mathrm{UNIONS5000}_{ugriz}$ but should not be trusted. We also provide $\mathtt{z\_ref}$ for the preferred redshift, and the SDSS spectroscopic redshifts are used if available; otherwise $\mathtt{Z\_B\_ref}$ is used.
@@
We calculate the spectral luminosities $\mathrm{\ W\ Hz^{-1}}$] at 1.4 and 3 GHz using the following equation:
@@
where $S_{\mathrm{obs}}$ [mJy] is the integrated flux density, $D_{L}$ is the luminosity distance, $z$ is the redshift, and $\alpha$ is the spectral index defined in the positive convention $S_{\mathrm{obs}}\propto\nu^{\alpha}$. The redshift adopts those stored in $\mathtt{`Z\_B\_ref^{\prime}}$ representing the UNIONS photometric redshifts. For ${L}_{\mathrm{3GHz}}$, if the UNIONS-VLASS source is detected in FIRST, then $\mathtt{alpha\_1400\_3000}$ is used, otherwise ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}$ is assumed to be 0.7-0.7 (condon_radio_1992; dermer_high_2009). For ${L}_{\mathrm{1.4GHz}}$, we adopt the following strategy for a UNIONS-VLASS source if it is:
  • detected in FIRST but not in LoTSS: $S_{\mathrm{obs}}=S_{\mathrm{FIRST}}$ and ${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}=-0.7$ used (gasperin_radio_2018);
@@
[…]
Figure 9: Left: Radio number counts normalized to the static Euclidean geometry. Right: VLASS 3 GHz source counts as a function of the AB radio magnitude t. Source counts at $11.5<t<15.5$ are fitted with a linear function, and the best fits are indicated by the colored line, while the solid ones are fittings in this work and dashed ones are from the literature. The black circles represent UNIONS-VLASS sources, black squares represent the “all” VLASS Epoch 2 catalog with criteria (except $\mathtt{NN\_dist}>30\arcsec$) in § 2.2 applied, and golden crosses represent UNIONS-VLASS sources detected in FIRST, and the size of each sample is shown in the parenthesis. Poisson errors in the sampling bins are represented by vertical bars.
@@ 4 Radio and Optical Number Counts
We here summarize the number counts and basic radio properties of UNIONS-VLASS radio galaxies, with comparisons from previous radio galaxy catalogs in the literature.
@@ 4.1 UNIONS-VLASS 3 GHz Radio Source Counts
Radio source number counts are often normalized to Euclidean geometry and examined on the plane of flux density ($S$) and $S^{2.5}\ dN/dS$, where $N$ is the number of sources within each flux density bin (left panel of Figure 9). Compared to the VLA-COSMOS 3 GHz Survey that reaches a median sensitivity of 2.3 $\mathrm{\mu Jy\ beam^{-1}}$ (smolcic_vla-cosmos_2017; indicated by violet diamonds), VLASS is limited in probing radio sources with ${F}_{\mathrm{int}}<1$ mJy. As a result, VLASS is comparatively biased toward brighter sources and is therefore more likely to select RL populations, which will be further discussed in § 4.3.
@@
We use also the plane of the radio AB magnitude t and the number counts N in each magnitude bin to make direct comparisons with FIRST catalogs at 1.4 GHz. Here, the radio AB magnitude is defined as:
[…]
@@
The $3\sigma\ (=0.36$ mJy) detection limit of VLASS then corresponds to 17.5 and the $5\sigma\ (=1$ mJy) detection limit of FIRST then corresponds to 16.4. In the right panel of Figure 9, we show radio source counts for 146,212 “clean” UNIONS-VLASS radio galaxies (black circles). We make a UNIONS-VLASS subsample for 79,638 sources that are also detected in FIRST (golden crosses) and compare it with 1.4 GHz radio galaxy catalogs employing FIRST (ivezic_optical_2002; yamashita_wide_2018). We also apply the criteria described in § 2.2, except $\mathtt{NN\_dist>30\arcsec}$, to the entire VLASS2 catalog, resulting in 1,873,519 sources (black squares). We compare this “all” VLASS2 sample with the “clean” VLASS2 sample that contains 1,380,707 sources used for the cross-match. The distributions of these samples at $11.5<t<15.5$ are fitted with a linear function and are indicated by the solid lines, and the best-fit parameters are summarized in Table 2. The best fits of the data collected from the literature are indicated by the dashed lines.
@@
The trends between “all” VLASS (blue solid line) and UNIONS-VLASS sources (black solid line) are in rough consistency. The global difference and the larger discrepancy at the bright end are attributed to the fact that we have applied $\mathtt{NN\_dist>30\arcsec}$ to VLASS2 in the pre-cross-match stage. This criterion may remove a significant fraction of low-$z$ sources that have bright radio lobes and hotspots identified as adjacent components, contributing to the number counts at the bright end. If we compare UNIONS-VLASS with the “clean” VLASS2 used for the cross-match, then this discrepancy diminishes, and the global difference is consistent with the matching rate of $0.594$ (left panel of Figure 9).
@@
Compared with 1.4 GHz HSC- and SDSS-FIRST source counts (golden and cyan dashed lines), our 3 GHz UNIONS-VLASS sources (black circles) have a slightly steeper increase at $11.5<t<15.5$. This could be explained by our choice of $\mathtt{NN\_dist>30\arcsec}$. The best fit of the subsample of our UNIONS-VLASS sources with FIRST detections (golden solid line) has a slope in general consistency with previous studies. This comes from the fact that the sky coverage of UNIONS-VLASS is beyond the survey area of FIRST, thus resulting in an underestimate of the real number counts for our UNIONS-VLASS-FIRST sample.
@@
All VLASS 3.509(±0.269)-3.509\ ($\pm$ 0.269) 0.317(±0.018)0.317\ ($\pm$ 0.018)
UNIONS-VLASS 4.870(±0.067)-4.870\ ($\pm$ 0.067) 0.391(±0.005)0.391\ ($\pm$ 0.005)
UNIONS-VLASS-FIRST 4.779(±0.142)-4.779\ ($\pm$ 0.142) 0.371(±0.010)0.371\ ($\pm$ 0.010)
$r$-band 0.932(±0.067)-0.932\ ($\pm$ 0.067) 0.074(±0.003)0.074\ ($\pm$ 0.003)